OpenStudy (tylerd):

A ball of radius 17 has a round hole of radius 8 drilled through its center. Find the volume of the resulting solid.

OpenStudy (agl202):

lid = 4π17^3 /3 - π(2*15)*8^2 - 2π2(3*8^2 + 2^2)/6 = 4π/3 * [4913 - 1440 - 98] = 3375 * 4π/3 = 4500π = 14137.167 cubic whatever unit. integration is the way asker's teacher/prof might require, but here i did using spherical cap's formula taken from wikipedia. of course it was derived using integration. i just took the easiest way, taking in all the processed material. asker never specified the approach he required. Hope this helped ya! =)

OpenStudy (tylerd):

this works for now.

OpenStudy (tylerd):

think you can help with 1 more?